OpenStudy (anonymous):

Use the distributive property to rewrite the expression and solve. fill in the blanks 6*94 = 6*?+ 6*4 = ?

OpenStudy (skullpatrol):

any ideas?

OpenStudy (anonymous):

not really

OpenStudy (skullpatrol):

ab + ac = a(b+c)

OpenStudy (anonymous):

k

OpenStudy (skullpatrol):

6*94 = 6*?+ 6*4 ______= 6( ?+ 4)

OpenStudy (anonymous):

k i think i see what your saying

OpenStudy (skullpatrol):

what goes here "?"

OpenStudy (anonymous):

94?

OpenStudy (skullpatrol):